OpenStudy (anonymous):

4/x-1+7/x+1=5x/x^2-1

OpenStudy (anonymous):

what is the Que ?

OpenStudy (anonymous):

4/x-1 + 7/x+1 = 5x/x^2-1 (4(x+1)+7(x-1))/x^2-1 = 5x/x^2-1 common denominators 4x+4+7x-7=5x denominators on both sides of the equation are same so can becancelled 11x-3=5x -3=5x-11x -3=-6x -3/-6 = x x=1/2